The solution to the system y = x(squared) + 2 and y = 2 consists of how many ordered pairs?
Found 2 solutions by ysl51, jpankey:
Answer by ysl51(1)
Answer by jpankey(3)

Substitute y=2 into the first equation, y = x^2 + 2.
2 = x^2 + 2, subtract 2 from both sides
0 = x^2, solve for x by taking square root of both sides
0 = x
There is one ordered pair (0,2) in this system